OpenStudy (anonymous):

What is the inverse function of y=4*(9)^x

OpenStudy (anonymous):

That y is equal to four times nine to the power of x

OpenStudy (freckles):

Have you tried solving for x yet?

OpenStudy (freckles):

you could first put it in this form: \[p=a^x \text{ then rewrite as } \log_a(p)=x\]

OpenStudy (freckles):

that is you divide both sides by 4 first you will have that desired form I mentioned .

OpenStudy (anonymous):

I am lost I don't know how to do this

OpenStudy (anonymous):

yo freckles log is too advanced original question is y=4*(9)^x inverse, flip the x and y x = 4*(9)^y solve for x

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sorry i meant solve for y

OpenStudy (anonymous):

I don't know how

OpenStudy (anonymous):

That link shows fractions my problem doesn't have fraction

OpenStudy (freckles):

\[\text{ In general, for something like: } y=b(a)^x \\ \text{ You can solve for } x \text{ by first doing } \\ \text{ division of } b \text{ on both sides } \\ \frac{y}{b}=a^x \\ \text{ now recall } \frac{y}{b}=a^x \text{ is equivalent to } \log_a(\frac{y}{b})=x\]

OpenStudy (freckles):

just using what I had above and replacing my p with y/b

OpenStudy (anonymous):

?